OpenStudy (baileedoherty):

A 6-foot-tall climber casts a 20-foot-long shadow at the same time that a tree casts a 90-foot-long shadow. How tall is the tree?

sam (.sam.):

Proportions right?

sam (.sam.):

If the climber is 6 foot tall and casts a shadow of 20 foot long, the ratio will be 6 : 20, so how tall the tree will be if it casts a 90 foot long shadow? Ratio for tree is ? : 90 We set it up like this \[6:20=x:90 \\ \\ \frac{6}{20}=\frac{x}{90} \\ \\ \frac{540}{20}=x \\ \\ x=?\]

OpenStudy (baileedoherty):

27?